# Hyperuniformity and optimal transport of point processes

### v3

@@ 3.3 Local to global allocation
Consider test functions $f,g\in C_{c}(\mathbb{R}^{d})$ supported by resp. $K,L$ compact regular sets. Let $\Omega_{n}$ be the event that $K-U_{n},L-U_{n}$ don’t hit $\partial\Lambda_{n}.$ There is a deterministic function $\theta$ such that, with $\theta_{K}:=\theta(K,U_{n}),\theta_{L}:=\theta(L,U_{n}),$ under $\Omega_{n}$
$(M_{n}(f,\mathbb{R}^{d}),M_{n}(\mathbb{R}^{d},g))$ $=(\tilde{\mu}_{n}(\tilde{f}),\tilde{\nu}_{n}(\tilde{g}))$
where $\tilde{f}(x)=f(x+\theta_{K}),\tilde{g}(x)=g(x+\theta_{L})$. The random variables $\tilde{\mu}_{n}(\tilde{f})$ and $\tilde{\nu}_{n}(\tilde{g})$ are independent because $\mu(\Lambda_{n}),\nu(\Lambda_{n})\textrm{ are independent and }{\mu_{n}}(\tilde{f}),{\nu_{n}}(\tilde{g})\textrm{ are independent.}$. Only the second point requires a proof: for Borel $I,J\subset\mathbb{R},$
$\mathbf{P}(\mu_{n}(\tilde{f})\in I,\nu_{n}(\tilde{g})\in J)=$ $\mathbf{E}(\mathbf{P}(\mu_{n}(\tilde{f})\in I,\nu_{n}(\tilde{g})\in J\;\mid\;U_{n}))$
$=$ ${\mathbf{E}(\mathbf{P}(\mu_{n}(\tilde{f})\in I\;\mid\;U_{n})\mathbf{P}(\nu_{n}(\tilde{g})\in J\;|\;U_{n}))}$
$=$ $\mathbf{P}(\mu_{n}(f)\in I)\mathbf{P}(\nu_{n}(g)\in J)$
where we have used that $\mu(\tilde{f}),\nu(\tilde{g})$ are independent conditionally on $U_{n}$ in the second equality and in the third equality, we use stationarity and independence with $U_{n}$ as well as independence between $\mu$ and $\nu.$
Note that $\tilde{\mu}_{n},\tilde{\nu}_{n}$ converge vaguely in law to $\mu,\nu$ respectively. So for $I,J$ open intervals of $\mathbb{R}$,
$\mathbf{P}(\mu(f)\in I)\mathbf{P}(\nu(g)\in J)=$ $\mathbf{P}(\mu(\tilde{f})\in I)\mathbf{P}(\nu(\tilde{g})\in J)$
$=$ $\lim_{n}\mathbf{P}(\tilde{\mu}_{n}(\tilde{f})\in I)\mathbf{P}(\tilde{\nu}_{n}(\tilde{g})\in J)$
$=$ $\lim_{n}\mathbf{P}(\tilde{\mu}_{n}(\tilde{f})\in I,\tilde{\nu}_{n}(\tilde{g})\in J)$
$=$ $\lim_{n}\mathbf{E}\Big{(}\mathbf{1}[\tilde{\mu}_{n}(\tilde{f})\in I,\tilde{\nu}_{n}(\tilde{g})\in J]\mathbf{1}[\Omega_{n}]\Big{)}+O\Big{(}\mathbf{P}(\Omega_{n}^{c})\Big{)}$
$=$ $\lim_{n}\mathbf{E}\Big{(}\mathbf{1}[M_{n}(f,\mathbb{R}^{d})\in I,M_{n}(\mathbb{R}^{d},g)\in J]\mathbf{1}[\Omega_{n}]\Big{)}+O(\mathbf{P}\Big{(}\Omega_{n}^{c})\Big{)}$
$=$ $\mathbf{P}(M(f,\mathbb{R}^{d})\in I,M(\mathbb{R}^{d},g)\in J)$
using $\mathbf{P}(\Omega_{n})\to 1$ and without loss of generality, assuming $M_{n}$ converge vaguely to $M$ in law. It means that $M$ is a coupling between an independent copy of $\mu$ and an independent copy of $\nu.$
We list the steps involved in the proof and only the second step needs a proof.
  1. (i)
    The overall strategy is to exhibit a $\mathbb{Z}^{d}$-stationary matching $T:\mu^{\prime}\to\mathbb{Z}^{d}$ where $\mu^{\prime}$ is distributed as $\mu$. Similarly, a $\mathbb{Z}^{d}$-stationary matching $T^{\prime}:\nu^{\prime}\to\mathbb{Z}^{d}$ also exists for $\nu^{\prime}$ distributed as $\nu$ and independent of $\mu^{\prime}$. We can shift the lattice by two uniform independent random variables to obtain $\mathbb{R}^{d}$-stationary matchings $\tilde{T},\tilde{T}^{\prime}$ and then $\tilde{T}^{\prime}\circ\tilde{T}^{-1}$ is a stationary matching between $\mu^{\prime}$ and $\nu^{\prime}$, with both being independent copies of $\mu$ and $\nu$ respectively. We now list the steps to produce a $\mathbb{Z}^{d}$-stationary matching.
  2. (ii)
    Using assumption on $\mu$ and the triangular inequality, we will show that $\tilde{\mu}_{n}$ has linear transport cost for $\bf Z_{n}$, ( i.e., $\mathbf{E}\widetilde{\mathsf{C}}_{w}(\tilde{\mu}_{n},\bf Z_{n})\leqslant cn$) where $\bf Z_{n}$ is obtained from $\mathbb{Z}^{d}\cap\Lambda_{n}$ by a slight rescaling and removing some points to match the number of atoms of $\tilde{\mu}_{n}$ such that the fraction of removed points goes to 00, hence $\bf Z_{n}\to\mathbb{Z}^{d}$ a.s. in the vague topology.
  3. (iii)
    Hence there is a coupling $m_{n}$ with linear cost between $\tilde{\mu}_{n}$ and $\bf Z_{n}$. Since $\tilde{\mu}_{n}$ and $\bf Z_{n}$ have same number of atoms with mass $n/N$, using Proposition 12, we have that the left hand side is realised by a transport map $T_{n}:\textrm{supp}(\tilde{\mu}_{n})\to\textrm{supp}(\bf Z_{n})$.
  4. (iv)
    Hence by Item (1) of the proposition, there is a distributional vague limit $M$ which has finite average cost. By Skohokod’s embedding theorem, $\tilde{\mu}_{n},{\bf Z_{n}},M_{n},M$ can be embedded in the same probability space so that the convergence is a.s. As in the third step of Part (1), we can then prove that $M$ is a coupling between a copy of $\mu$ and $\mathbb{Z}^{d}$, or said differently, $\mu^{\prime}:=M(\cdot,\mathbb{R}^{d})$ and $Z^{\prime}:=M(\mathbb{R}^{d},\cdot)$ are independent and distributed as $\mu,\mathbb{Z}^{d}$ respectively. Since a.s., ${\bf Z_{n}},\tilde{\mu}_{n}$ converge vaguely to $Z^{\prime},\mu^{\prime}$ respectively, by Lemma 13(2) $M$ gives a matching between $\mu^{\prime}$ and $Z^{\prime}$.
We need to only prove the second step (Step (ii)) i.e., construction of ${\bf Z_{n}}$ with linear transport cost between $\tilde{\mu}_{n}$ and $\bf Z_{n}$.
Construction of ${\bf Z_{n}}$ with linear transport cost and convergence ${\bf Z_{n}\to\mathbb{Z}^{d}}$:
Let $N=\mu(\Lambda_{n}),\mu_{n}=\mu 1_{\Lambda_{n}},\tilde{\mu}_{n}=\frac{n}{N}\mu_{n}$. By integrability of the RPCM $\beta$, $n^{-2}\textrm{Var}(N)\to 0$ and through Bienaymé-Chebyshev’s inequality, we obtain $n^{-1}N\to 1$ in probability. Thus we can choose a subsequence that converges a.s. and we will work with this subsequence now onwards. Up to change the definition of $\Lambda_{n}$, we can assume $n^{-1}N\to 1$ a.s..
The strategy is to show that $\tilde{\mu}_{n}$ is close to $A_{n}\mathbb{Z}^{d}\cap\Lambda_{n}$ for some random sequence $A_{n}\to 1$, using that $A_{n}\mathbb{Z}^{d}\cap\Lambda_{n}\to\mathbb{Z}^{d}$ a.s. in the vague convergence. We wish to choose the scaling $A_{n}$ so that the number of points match: #$\#A_{n}\mathbb{Z}^{d}\cap\Lambda_{n}=N$, but unfortunately for $a>0$, #$\#(a\mathbb{Z}^{d})\cap\Lambda_{n}$ is of the form $m^{d}$ for some $m\in\mathbb{N}$, and $N$ might not be of this form, which is a source of nuisance. Let $m\in\mathbb{N}$ so that $(m-1)^{d}<N\leqslant m^{d}$. Let $A_{n}^{d}=n/m^{d}>0$ so that
#$\#(A_{n}\mathbb{Z}^{d})\cap\Lambda_{n}=m^{d}=N+p\textrm{ where }p\leqslant m^{d}-(m-1)^{d}\leqslant dm^{d-1}.$
Since $N\to\infty$ a.s., we have that $m\to\infty$ a.s. and further by the above bounds, $m^{-d}p\leq dm^{-1}\to 0$ a.s.. This yields that $N/m^{d}\to 1$ a.s. and so $A_{n}=n/m^{d}\to 1$ a.s. as well.
If $p=0$ (i.e., $N=m^{d}$), set
${\bf Z}_{n}:=\sum_{k\in A_{n}\mathbb{Z}^{d}\cap\Lambda_{n}}\frac{n}{N}\delta_{k}.$
Otherwise, we build ${\bf Z_{n}}$ by removing $p$ points carefully from $A_{n}\mathbb{Z}^{d}\cap\Lambda_{n}$, so that ${\bf Z_{n}}$ and $\tilde{\mu}_{n}$ both have $N$ atoms, each of which is assigned mass $n/N$, and so that the transport cost stays under control. Formally,
${\bf Z}_{n}:=\sum_{k\in A_{n}\mathbb{Z}^{d}\cap\Lambda_{n}\setminus B_{n}}\frac{n}{N}\delta_{k}$
where we define the set of removed points $B_{n}$ below, leading to $|B_{n}|=p$. The proportion of removed points is bounded by $(m^{d}-N)/m^{d}$, hence it goes to 00, and we prove below that ${\bf Z_{n}}\to\mathbb{Z}^{d}$ a.s. in the vague topology.
We now elaborate on the procedure of removing the points so that ${\bf Z_{n}}$ also has a nice transport plan with ~$\tilde{\mathcal{L}}_{n}$ (and hence further to $\tilde{\mu}_{n}$). Ultimately each of the $N$ points of ${\bf Z_{n}}$ must be allocated to an area of volume $\frac{n}{N}$. The idea is to partition $\Lambda_{n}$ into $p$ nearly-equal blocks, with each block consisting of approximately $\frac{m^{d}}{p}=\frac{N}{p}+1$ boxes centred at points of $A_{n}\mathbb{Z}^{d}\cap\Lambda_{n}$. Then in each block, we remove one of the points and re-allocate ‘efficiently’ the mass (approximately $\frac{m^{d}}{p}A_{n}^{d}=\frac{n}{p}$) to the remaining points ($\frac{N}{p}$) and thereby giving each point a mass of $\frac{n}{N}$ as required.
Let us number $z_{1},\dots,z_{m^{d}}$ the points of $A_{n}\mathbb{Z}^{d}\cap\Lambda_{n}$ so that $z_{i}$ is a neighbour of $z_{i+1}$ (i.e. at distance $A_{n}$). We allocate to $z_{1}$, the mass from the surrounding cube of measure $\frac{n}{m^{d}}\leqslant\frac{n}{N}$ around $z_{1}$, but there is still a missing unallocated mass $\delta_{n}:=\frac{n}{N}-\frac{n}{m^{d}}=\frac{np}{Nm^{d}}>0$. We borrow this defect mass $\delta_{n}$ from the next cube i.e., box of $z_{2}$, in an arbitrary way, which we represent by an arrow going from $z_{2}$ to $z_{1}$. Now we allocate the remaining mass $\frac{n}{m^{d}}-\delta_{n}$ in the box of $z_{2}$ to itself and borrow remaining mass $2\delta_{n}$ from the box of $z_{3}$, i.e. we put two arrows from $z_{3}$ to $z_{2}$. We do this successively. After allocating $k$ points, $z_{k}$ will have to borrow $k\delta_{n}$ mass from the box of $z_{k+1}$, represented by $k$ arrows. Let ${k_{1}}$ be the last $k$ such that $k\delta_{n}\leqslant\frac{n}{m^{d}}$. We remove the $(k_{1}+1)$-th point from $A_{n}\mathbb{Z}^{d}\cap\Lambda_{n}$ or equivalently retain $z_{1},\ldots,z_{k_{1}}$ to $\bf Z_{n}$. This is called the first batch. Note that $k_{1}\leqslant N/p$ and that each $z_{i}$ borrows mass from its own box or the neighbouring box of $z_{i+1}$. So mass is transported at most distance of $2A_{n}$, i.e. each arrow is long of at most $2A_{n}$, and the total mass carried from a point to its neighbour is smaller or equal to $n/N$, the total transport cost of this first step is hence bounded by
${\sum_{i=1}^{k_{0}}\underbrace{w(2A_{n})}_{\textrm{ distance }}\underbrace{\frac{n}{N}}_{\textrm{ mass }}\leqslant c\,\frac{k_{1}nA_{n}^{d}}{N}\leqslant c\,\frac{n\frac{n}{m^{d}}}{N}\frac{n}{m^{d}}\frac{Nm^{d}}{np}\leqslant c\,\frac{n^{2}}{pm^{d}}}.$
This procedure is repeated at most $p=N-m^{d}$ times. Each time, the $j$-th batch $\llbracket b_{j-1},b_{j}\rrbracket:=\{b_{j-1},\dots,b_{j}\},$ with $b_{0}=0$, ouputs an excess mass denoted by $\delta_{n,j}\in[0,\delta_{n}]$, and this mass is allocated to the first point in the $(j+1)$-th batch (with the convention $\delta_{n,0}=0$). Then we follow a similar procedure: each point of the $(j+1)-$th batch borrows from its right neighbours $\delta_{n}$ more mass than it gives to its left neighbour, and we denote by $k_{j}$ the last integer such that the total mass borrowed from the right neighbour is $\leqslant n/N$. Note that $k_{j}\in\{k_{1}-1,k_{1}\}$. This leaves a residue mass $\delta_{n,j+1}\in[0,\delta_{n}]$. The same reasoning yields that the cost for each step is upper bounded by $\frac{n^{2}}{pm^{d}}$ and so the total cost is bounded by
${C_{w}({\bf Z_{n}},\tilde{\mathcal{L}}_{n})\leqslant c\,\frac{n^{2}}{m^{d}}.}$
Heuristically, we expect $m^{d}\sim n$ and thus the above bound yields $C_{w}({\bf Z_{n}},\tilde{\mathcal{L}}_{n})\leqslant n$, thereby completing the proof. More formally, if we show that $\mathbf{E}[\frac{1}{N}\mathbf{1}[N\geq 1]]\leqslant cn^{-1}$, we can derive that
$\mathbf{E}[C_{w}({\bf Z_{n}},\tilde{\mathcal{L}}_{n})\mathbf{1}[N\geq 1]]\leqslant n^{2}\mathbf{E}[\frac{1}{m^{d}}\mathbf{1}[N\geq 1]]\leqslant n^{2}\mathbf{E}[\frac{1}{N}\mathbf{1}[N\geq 1]]\leq cn.$
Using Bienaymé-Chebyshev’s inequality, we can derive the requisite bound as follows.
$\mathbf{E}[\frac{1}{N}\mathbf{1}[N\geq 1]]$ $\leqslant cn^{-1}+c\mathbf{P}(1\leqslant N<n/2)\leqslant cn^{-1}+cn^{-2}\textrm{Var}(N)\leqslant cn^{-1}.$
The above transport cost bound combined with the assumption on $\tilde{\mu}_{n}$ and triangle inequality, gives us that $\mathbf{E}[C_{w}({\bf Z_{n}},\tilde{\mu}_{n})\mathbf{1}[N\geq 1]]\leqslant cn$. In case $N=0$, $C_{w}({\bf Z_{n}},\tilde{\mu}_{n})=0$ trivially and therefore the claim in Step (ii) above follows.
Before proceeding further, we must actually modify slightly the previous construction, without increasing the transport cost, i.e. the number of arrows carrying mass at most $n/N$ and going from one point to one of its neighbour, the procedure is illustrated by Figure 1:
  • instead of removing the point with index $b_{j}$ in the $j$-th batch, we remove a point uniformly at random with an arbitrary index $I_{j}\in\llbracket b_{j-1}+1,b_{j}\rrbracket$. We do this independently for each of the $p$ blocks.
  • We reorganise the “borrowing arrows” of the previous construction, so that the removed point has $1$ outgoing arrow of mass $\delta_{n}-\delta_{n,j}$ towards its left neighbour, one outgoing arrow of mass $\delta_{n,j+1}$ towards its right neighbour, and $k_{j}-2$ outgoing arrows of mass $\delta_{n}$, and each other point has not more than a total $n/N$ mass going in or out. Note that the $\delta_{n,j+1}$ mass will be successively passed along from point $I_{j}$ until the first point of the next batch, but, from the very definition of $k_{j}$, the total mass going from each box to a neighbour box will not exceed a maximal value of $n/N$.
$\delta_{n,j}$$\delta_{n}-\delta_{n,j}$$\delta_{n,j+1}$$\delta_{n,j+1}$$\delta_{n}$$\delta_{n}$$\delta_{n}$$\delta_{n}$$\delta_{n}$$\delta_{n}$$b_{j-1}+1$$I_{j}$$b_{j}$
Figure 1: reallocation of mass: the total $n/N$ mass of the $I_{j}$’th box is reallocated to other boxes of the $j$-th batch in such a way that (i) each other box receives $\delta_{n}$ more mass than it gives, (ii) the transported mass always goes from one box to a neighbour box (where it might be transported again), (iii) the mass transported from a box to a neighbour box is bounded by $n/N$.
The number of arrows and their masses do not increase (in fact, they will strictly decrease in general if $I_{j}$ is more “towards the middle of the batch”), the bound on the transport cost is still valid. We explain now how we choose $I_{j}$ (and why). The goal of this modification is to ensure the a.s. vague convergence ${\bf Z}_{n}\to\mathbb{Z}^{d}$. We know that since $A_{n}\to 1$ a.s., $A_{n}\mathbb{Z}^{d}\cap\Lambda_{n}\to\mathbb{Z}^{d}$ in the vague convergence a.s., i.e. for any $A$ without integer point on the boundary, #$\#A_{n}\mathbb{Z}^{d}\cap\Lambda_{n}\cap A\to\#\mathbb{Z}^{d}\cap A$. So it simply remains to prove that for any such $A$ without integer point on the boundary, #$\#A_{n}\mathbb{Z}^{d}\cap\Lambda_{n}\cap A-\#{\bf Z}_{n}\cap A\to 0$ a.s. This amounts to proving that for $K\in\mathbb{N}$, for some random $n_{K}$ and $n\geqslant n_{K}$, there are a.s. no removed points in $B(0,K)$. The density of removed points goes to 00 as $p/N\to 0$, but we must ensure that
@@
Up to further taking a subsequence, this will ensure the conclusion that for $n\geqslant n_{K}$, a.s. all the points in $B(0,K)$ remain. So that it works simultanously for all $K$, we must simply choose $n_{K+1}\geqslant n_{K}$ a.s.. So that (43) is in order, we decided to choose $I_{j}$ randomly, independently and uniformly in $\llbracket b_{j-1}+1,b_{j}\rrbracket$ in the previous modified construction. Each $z_{k},k\in Z_{j}$ has a (conditional) probability $\frac{1}{b_{j}-b_{j-1}}=k_{j}^{-1}$ to be removed. On the event $\Omega_{n}=\{|N-n|\leqslant n/2\}$, whose probability goes to $1$ with Bienaymé-Chebyshev inequality, we therefore have #$\#A_{n}\mathbb{Z}^{d}\cap K\leqslant C_{K}<\infty$. We also have conditionally on $\Omega_{n}$, since by definition $(k_{j}+1)\delta_{n}\geqslant 1$, and
$\delta_{n}=2(\frac{n}{N}-\frac{n}{m^{d}})\to 0,$
$\max_{j}k_{j}\to\infty$ and the union bound yields
$\mathbf{P}(\text{\rm{there is a removed point in $B(0,K)$}}\mid\Omega_{n})\leqslant C_{K}\max_{k}\mathbf{P}(z_{k}\text{\rm{ removed}}\ \mid\,\Omega_{n})\to 0,$
which concludes the proof.

### v4

@@ 3.3 Local to global allocation
Consider non-negative test functions $f,g\in C_{c}(\mathbb{R}^{d})$ supported by resp. $K,L$ compact convex sets. For $u\in\mathbb{R}^{d}$, let $\theta(K,u)$ be $\infty$ if KuK-u touches one of the boundaries $\partial(z2\pi n^{1/d}+\Lambda_{n}),z\in\mathbb{Z}^{d}$, and otherwise it is the element $z\in\mathbb{Z}^{d}$ such that $K-u\subset(2\pi n^{1/d}z+\Lambda_{n})$. Hence, under the event $\Omega_{n}:=\{\theta(K,U_{n})\neq\infty\}$
$(M_{n}(f,\mathbb{R}^{d}),M_{n}(\mathbb{R}^{d},g))$ $)\displaystyle=(\tilde{\mu}_{n}(\tilde{f}),\tilde{\nu}_{n}(\tilde{g}$))
@@
$\mathbf{E}e^{-\mu_{n}(\tilde{f})-\nu_{n}(\tilde{g})}$ $=\mathbf{E}\bigl(e^{-\mu_{n}(\tilde{f})-\nu_{n}(\tilde{g})}\;\mid\;U_{n})\bigr)$
$={\mathbf{E}\bigl(\mathbf{E}(e^{-\mu_{n}(\tilde{f})}\;\mid\;U_{n})\,\mathbf{E}(e^{-\nu_{n}(\tilde{g})}\;|\;U_{n})\bigr)}$
[…]
$\mathbf{E}e^{-\mu(f)}\mathbf{E}e^{-\nu(g)}=$ $\mathbf{E}e^{-\mu(\tilde{f})}\mathbf{E}e^{-\nu(\tilde{g})}$
$=$ $\lim_{n}\mathbf{E}e^{-\tilde{\mu}_{n}(\tilde{f})}\,\mathbf{E}e^{-\tilde{\nu}_{n}(\tilde{g})}$
$=$ $\lim_{n}\mathbf{E}e^{-\tilde{\mu}_{n}(\tilde{f})-\tilde{\nu}_{n}(\tilde{g})}$
$=$ $\lim_{n}\mathbf{E}\Big(e^{-\tilde{\mu}_{n}(\tilde{f})-\tilde{\nu}_{n}(\tilde{g})}\mathbf{1}_{\Omega_{n}}\Big)+O\Big(\mathbf{P}(\Omega_{n}^{c})\Big)$
$=$ $\lim_{n}\mathbf{E}\Big(e^{-M_{n}(f,\mathbb{R}^{d})-M_{n}(\mathbb{R}^{d},g)}\mathbf{1}_{\Omega_{n}}\Big)+O(\mathbf{P}\Big(\Omega_{n}^{c})\Big)$
$=$ $\mathbf{E}e^{-M(f,\mathbb{R}^{d})-M(\mathbb{R}^{d},g)}$
using $\mathbf{P}(\Omega_{n})\to 1$ and without loss of generality, assuming $M_{n}$ converge vaguely to $M$ in law. Thus we have that
$\mathbf{E}e^{-(M(f,\mathbb{R}^{d})+M(\mathbb{R}^{d},g))}=\mathbf{E}e^{-\mu(f)}\,\mathbf{E}e^{-\nu(g)}=\mathbf{E}e^{-M(f,\mathbb{R}^{d})}\,\mathbf{E}e^{-M(\mathbb{R}^{d},g)}.$
Now given compact sets $K,L$ and $s,t\geq 0$, there exist sequence of bounded continuous functions $f_{n},g_{n}$ such that $f_{n}\downarrow t\mathbf{1}_{K}$ and $g_{n}\downarrow s\mathbf{1}_{L}$ respectively. By monotonic approximation, the above identity yields that
$\mathbf{E}e^{-tM(K,\mathbb{R}^{d})-sM(\mathbb{R}^{d},L)}=\mathbf{E}e^{-tM(K,\mathbb{R}^{d})}\,\mathbf{E}e^{-sM(\mathbb{R}^{d},L)},$
thereby showing the independence of $(M(K,\mathbb{R}^{d}),M(\mathbb{R}^{d},L))$ for all compact sets $K,L$. This guarantees the independence of $(M(\cdot,\mathbb{R}^{d}),M(\mathbb{R}^{d},\cdot))$ via [6, Corollary 1.3.15] and that $M(\cdot,\mathbb{R}^{d}),M(\mathbb{R}^{d},\cdot)$ are simple point processes.
Fourth Step: if $\mu,\nu$ are simple independent point processes, $M$ can be chosen to be a.s. a matching
This step is actually similar to the proof of Proposition 2.11 in [25]. To be precise, they define a distance $\mathsf{C}(\mathscr{L}_{\mu},\mathscr{L}_{\nu})$ between the laws $\mathscr{L}_{\mu},\mathscr{L}_{\nu}$ of resp. $\mu,\nu,$ which satisfies with their Proposition 2.4
$\mathsf{C}(\mathscr{L}_{\mu},\mathscr{L}_{\nu})\leqslant\inf_{M}\mathbf{E}\int_{\Lambda_{1}\times\mathbb{R}^{d}}w(x-y)M(dx,dy)$
where the infimum is over invariant couplings $M$ between independent versions $\mu^{\prime},\nu^{\prime}$ of $\mu$ and $\nu$ (note that independence is not required in their result). The RHS in fact corresponds with Fubini’s theorem to the definition of the typical distance from (3) for the equivariant matching $M$ built at first step above (the invariance comes from (47) which passes to the limit when $M_{n}\to M$) : by definition of the typical distance,
$\mathbf{E}w(X)=\int_{\mathbb{R}_{+}}\mathbf{E}\int{\bf 1}\{w(\|x-y\|)\geqslant r\}M(dx,dy)dr=\mathbf{E}\int\underbrace{\int{\bf 1}\{w(\|x-y\|)\geqslant r\}dr}_{=w(\|x-y\|)}M(dx,dy).$
Then [25, Proposition 2.11] allows to conclude directly that there is an equivariant matching between two versions $\mu^{\prime},\nu^{\prime}$ of $\mu,\nu$ which typical distance $X_{\text{\rm{match}}}$ satisfies $\mathbf{E}w(X_{\text{\rm{match}}})<\infty.$
Remark 4.
In the result of [25], it is not assumed that the point processes are simple. In fact, if they are not simple, there is not always a matching satisfying the conclusion of Proposition 11. It does not contradict [25, Proposition 2.11] because a more general definition of a matching is used. Specifically, they define matching as a point process in the product space but not necessarily induced by an actual bijection. However, we define matchings via bijections and hence we need the assumption of simplicity for point processes.